A round wastebasket with radiys 0. 15m opening in a uniForm electric field of 300N/C prependicular to opening.waht is the total electric flux through bottom and sides?
1
Expert's answer
2020-03-25T11:10:57-0400
"\Phi_{out}=\Phi_{in}=EA"

"\Phi_{out}=4\pi r^2E=4\pi (0.15)^2(300)=85\ Vm"